OpenStudy (anonymous):

Provide a 1-variable linear equation of your own creation. Explain the techniques, and show the steps used for solving the equation. Check that your solution is correct.

OpenStudy (anonymous):

First of all, what's a variable linear equation? "An algebraic equation, such as y = 2x + 7 or 3x + 2y − z = 4, in which the highest degree term in the variable or variables is of the first degree. The graph of such an equation is a straight line if there are two variables." Can you understand that?

OpenStudy (anonymous):

If so, then all you've got to do is to invent an equation with how many variables you like, but the highest degree term has to be 1.

OpenStudy (anonymous):

For instance: 3x - 2 = y, 2a + 3b = -4c.

OpenStudy (anonymous):

Oh, attention. Your exercise says it has to be only 1 variable.

OpenStudy (anonymous):

Like: x - 3 = 8 or 2y + 3 =1